OpenStudy (anonymous):

-100 + 5 + (-3) (1 point)

OpenStudy (anonymous):

-98

OpenStudy (anonymous):

-98

OpenStudy (karatechopper):

-98

OpenStudy (anonymous):

-100+5+(-3) = -100+5-3 = -103+5 = -98

OpenStudy (anonymous):

-98

OpenStudy (valpey):

You should give us more points